Janelle and Sandra are planting various colors of flowers in a circular shape on the front lawn of their university to display the school's new logo. The circle has a diameter of 12 feet and the flowers cost $15.00 per square yard. Find the cost of the flowers.
Answer by stanbon(75887)

Janelle and Sandra are planting various colors of flowers in a circular shape on the front lawn of their university to display the school's new logo. The circle has a diameter of 12 feet and the flowers cost $15.00 per square yard. Find the cost of the flowers.
----
diameter = 12 feet = 4 yds
radius = 2 yds
-------------------
Area of planted circle = pi*2^2 = 4pi sq. yrds
--------------------
Ans:: Cost = 15*4pi = 60pi dollars = $188.50
--------------
Cheers,
Stan H.
-----------------